EDITORIAL

Terry Goodman

This month I will be attending the Missouri Statewide Transfer and Articulation Conference where the focus will be on mathematics. Particular attention will be given to developing a more coherent set of expectations in mathematics for students in the last two years of high school and the first year of college. Indeed, the Coordinating Board for Higher Education and the Department of Elementary and Secondary Education have appointed a Mathematics Technical Advisory Committee to coordinate efforts in this area.

I believe that it is important for those of us in the mathematics community to be actively involved in this discussion. There are a number of important issues/questions to consider as we investigate ways to help students make a successful transition from high school to college level mathematics. Examples of such questions are:

1. How should college and universities accept/place students who come from high schools that have an integrated mathematics curriculum?

2. What criteria should colleges and universities use for placing students in the proper mathematics course?

3. Where do we find models of effective placement programs?

4. How can colleges and universities best use early placement analysis such as the Ohio State Early Math Placement Test?

5. How can college and university mathematics departments effectively communicate to high schools the mathematics they expect incoming students to know?

6. What mathematics do educated citizens need to know?

7. What are the crucial issues related to dual credit mathematics courses?

These and related questions will almost certainly have no simple answers. I would encourage you to provide input concerning such issues when you have the opportunity. We at MJMS would welcome any comments you have concerning these and other related questions.